40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 3249     Entry time: Tue Jul 20 11:49:31 2010
Author: Jenne 
Type: Update 
Category: SUS 
Subject: Some Suspensions not damping 

[Jenne, Koji]

I moseyed into the control room this morning, to find ITMX and ITMY both with their watchdogs tripped.  ITMY (new convention) wouldn't damp.  Koji discovered that there was a sign flip in 2 of the sensors.  A set of reboots of c1susvme1&2 fixed the problem. 

A side note:  For the ETMs, the OSEM sensor readouts are gigantic (~20,000), whereas for the similar channels on all other optics, the readouts are on the order of 1.  After some looking around, it seems that this is just the way things have been (for at least 100 days), and the filters in the SUSPOS and other SUS filter banks have a high pass filter to take care of this.  It's weird, but it seems to be the way it is, and the ETMs damp, so it's all good.

ELOG V3.1.3-